Serenity currently sells televisions for Walter's TV Store at a salary of $16,000 plus a $15 commission for each television she sells. Kristen's TV Store offers her a position with a salary of $10,400 plus a commission of $35 for each television she sells. How many televisions would she need to sell for the options to be equal?

Answer by ikleyn(52754)

.
Serenity currently sells televisions for Walter's TV Store at a salary of $16,000
plus a $15 commission for each television she sells.
Kristen's TV Store offers her a position with a salary of $10,400
plus a commission of $35 for each television she sells.
How many televisions would she need to sell for the options to be equal?
~~~~~~~~~~~~~~~~~~

Let t be the number of televisions under the problem's question.

According to the problem, we can write

    at Walter's  TV store the salary is  16000 + 15*t;

    at Kristen's TV store the salary is  10400 + 35*t.


The salaries are equal when

    16000 + 15*t = 10400 + 35*t.


Simplify this equation and find t from it

    16000 - 10400 = 35t - 15t

         5600     =    20t

            t     =   5600/20 = 280.


ANSWER.  The salaries are the same when 280 televisions are sold.

Solved.

---------------

Very standard problem with very standard solution.

Level of complexity is very basic.

Average student should solve such problems instantaneously without thinking.


        Come again to this forum soon to learn something new.